# On stable assignments generated by choice functions of mixed type

### v1

@@ 6 The poset of rotations
We start with analyzing some details of the ordinary method. Recall that a current, $i$-th, iteration in it consists of two stages: at the 1st one, the current boundary function $b^{i}$ is transformed into $x^{i}$, by setting $x^{i}_{f}:=C_{f}(b^{i}_{f})$ for each $f\in F$, and at the 2nd one, $x^{i}$ is transformed into $y^{i}$, by setting $y^{i}_{w}:=C_{w}(x^{i}_{w})$ for each $w\in W$. Then the boundary function is updated according to rule (3.1). Let $F^{i=}$ denote the set of fully filled vertices $f$ in $F$ for $x^{i}$ (i.e., $|x^{i}_{f}|=q(f)$), and $W^{i=}$ the set of fully filled vertices in $W$ for $y^{i}$. For $f\in F$, let $L^{i}_{f}$ denote the set of edges $e\in E_{f}$ such that $b^{i}(e)<b(e)$. We observe the following.
(i) For each $f\in F$, the sets $L_{f}$ are monotone non-decreasing: $L^{i}_{f}\subseteq L^{i+1}_{f}$, and the values on each edge $e\in L_{f}$ may decrease (at the 2nd stages of some iterations) but never increase (since $b^{1}(e)\geq\cdots\geq b^{i}(e)\geq b^{i+1}(e)\geq\cdots$).
@@
(iii) The heads $H_{w}(y)$ of vertices $w\in W^{\bullet=}$ are monotone non-decreasing: $H_{w}(y^{i})\subseteq H_{w}(y^{i+1})$ (when $w\in W^{i=}$), unless the critical tie in $\Pi_{w}$ changes (becomes better). In turn, the heads $H_{j}(x)$ for $f\in F$ are monotone non-increasing: $H_{f}(x^{i})\supseteq H_{f}(x^{i+1})$ (taking into account that $H_{f}(x^{j})=\emptyset$ if $f\in F-F^{j=}$).
@@
(iv) For each $f\in F$, all edges in $E_{f}-(H^{i}_{f}\cup L^{i}_{j})$ are saturated, where $H^{i}_{f}$ denotes the head $H_{f}(x^{i})$ (which is nonempty if $f\in F^{i=}$, and empty if $f\in F-F^{i=}$).
@@
These observations point out monotone parameters, and we say that $i$-th iteration is positive if at least one of the following events happens in it: for some $f\in F$, the set of edges $e\in E_{f}$ that are either saturated or contained in $L_{f}$ increases; the set $W^{\bullet=}$ of fully filled vertices in $W$ increases; for some $w\in W^{\bullet=}$, either the head $H_{w}$ increases, or the critical tie for $w$ becomes better. The number of positive iterations does not exceed $2|E|+|W|$.
@@
Next we deal with a sequence $Q$ of consecutive non-positive iterations and consider a current iteration among these. At the 1st stage of this iteration, the sum of values of the current function $x$ increases (compared with the preceding $y$) by some amount $\xi>0$. Note that if for a vertex $w\in W^{\bullet=}$, the value increases on some edge $e=fw$, then this edge cannot belong to the head $H_{w}$ or to a tie in $\Pi^{\rm end}_{w}$. (For otherwise the cutting operation at the 2nd stage of the current iteration decreases the value on $e$, and as a consequence, $e$ is added as a new element to the set $L_{f}$, which is impossible since the iteration is non-positive.)
@@
It follows that an increase on some edge $e=fw$ at the 1st stage of an iteration in $Q$ is possible only in two cases: either $w\in W^{\bullet=}$ and $e$ is contained in the current tail $T_{w}$, or $w$ belongs to the current deficit set $W-W^{\bullet=}$. This implies that the total increase $\xi^{\prime}$ on the edges incident to vertices in $W^{\bullet=}$ satisfies $\xi^{\prime}\leq\xi$ (the equality may happen only if there is no contribution from $\xi$ to the vertices in $W-W^{\bullet=}$). At the 2nd stage of this iteration, the total value in the heads $H_{w}$ for these vertices $w\in W^{\bullet=}$ must decrease by the same amount $\xi^{\prime}$ (emphasize that these heads $H_{w}$ do not change, and each edge $fw$ where the value decreases has been added to the set $L_{f}$ on earlier iterations). Then we come to the 1st stage of the next iteration with the total “quota-underloading” $\xi^{\prime}$ in corresponding vertices in $F$, which further causes a “quota-overloading” $\xi^{\prime\prime}\leq\xi^{\prime}$ in $W^{\bullet=}$, and so on. When such “overloadings” decrease geometrically, non-positive iterations may continue infinitely long (yet providing a convergence to a limit function).
@@
Relying on the observations above, we now are going to aggregate non-positive iterations from the ordinary method. More precisely, consider the first iteration in the sequence $Q$; assume that it is $i$-th iteration in the ordinary process. For brevity, we denote $x^{i}$ by $x$, and $y^{i}$ by $y$ (the functions obtained on the 1st and 2nd stages of the first iteration in $Q$). Also denote:
@@
Our idea consists in replacing the sequence $Q$ without the first iteration in it by one “big” (aggregated) iteration. To do so, we arrange the following linear program with variables $\phi_{f}$ for $f\in F^{=}$, and $\psi_{w}$ for $w\in W^{=}$:
@@
Define $y^{\prime}:=y+\Delta^{\phi,\psi}$. Then, in light of the properties of $y$, we observe the following: (A.2) implies that all vertices in $W^{=}$ are fully filled for $y^{\prime}$; (A.3) implies that $y^{\prime}$ is nonnegative; (A.4) implies that $y^{\prime}$ does not violate the quotas within $F^{=}$; (A.5) implies that $y^{\prime}$ is does not exceed the upper bounds $b$; (A.6) can be rewritten as $y^{\prime}(e)=a(y,w)-\psi_{w}\geq y(fw)+\phi_{f}=y^{\prime}(fw)$, where $e\in H_{w}$, implying that the edges in the head $H_{w}$ for $y$ belong to the head at $w$ for $y^{\prime}$ as well; and (A.7) implies that $y^{\prime}$ does not violate the quotas within $W-W^{=}$.
@@
Now assume that none of the constraints in (A.3),(A.5)–(A.7) turns into equality, and therefore, the attained constraints concern only (A.4). Then the resulting function $y^{\prime}$ is well-defined, in the sense that it is suitable to applying further ordinary iterations. Let $x^{\prime\prime}$ be the function obtained by applying to $y^{\prime}$ the 1st stage of the ordinary iteration. We observe that if ${\cal L}$ involves all inequalities in (A.4), then all vertices in $F^{=}$ are fully filled for $y^{\prime}$; this implies that $x^{\prime\prime}=y^{\prime}$, and the process terminates.
Now suppose that some inequalities in (A.4) are strict for $y^{\prime}$, and therefore, the subset $F^{\prime}$ formed by $f\in F^{=}$ with $|y^{\prime}_{f}|<q(f)$ is nonempty. Let $y^{\prime\prime}$ be the functions obtained by applying to $x^{\prime\prime}$ the 2nd stage of the ordinary iteration. Then $\delta:=y^{\prime\prime}-y^{\prime}$ is represented as $\Delta^{\phi^{\prime},\psi^{\prime}}$, where $\phi^{\prime}$ is a function on $F^{=}$ such that $\phi^{\prime}_{f}>0$ for each $f\in F^{\prime}$. Now form the pair $(\phi^{\prime\prime},\psi^{\prime\prime})=(\phi+\phi^{\prime},\psi+\psi^{\prime})$. One can see that this $(\phi^{\prime\prime},\psi^{\prime\prime})$ is a feasible solution to (A.1)–(A.8). But $\eta(\phi^{\prime\prime})=\eta(\phi)+\eta(\phi^{\prime})>\eta(\phi)$, contrary to the optimality of $(\phi,\psi)$.
Thus, the big iteration either terminates the whole process or changes at least one of the above parameters. The modified method that we elaborate is organized so that each big iteration should be preceded by an ordinary iteration and there are no two consecutive ordinary non-positive iterations. Therefore, it finishes after $O(|E|)$ iterations. Each big iteration is performed in time polynomial in $|E|$, by using a strongly polynomial l.p. method (see, e.g., [15]) to solve linear program (A.1)–(A.8).
@@
It remains to ask: whether or not the resulting $\widehat{x}$ found by the modified algorithm does coincide with the assignment $x^{\rm min}$? In essence this question is not so important to us, because once we are given one or another stable assignment, we can always transform it into $x^{\rm min}$ by applying a sequence of $O(|E|)$ rotational shifts, by considering the polar order $\succeq_{W}$ on ${\cal S}$ and acting as in Sect. 6. As a result, we can conclude with the following
@@
The modified algorithm is finite and finds the assignment $x^{\rm min}$. It consists of $O(|E|)$ iteration, each taking time polynomial in $|E|$. normal-$\Box$

### v2

@@ 7 Affine representability and stable assignments of minimum cost
We start with analyzing some details of the ordinary method. Recall that a current, $i$-th, iteration in it consists of two stages: at the 1st one, the current boundary function $b^{i}$ is transformed into $x^{i}$, by setting $x^{i}_{f}:=C_{f}(b^{i}_{f})$ for each $f\in F$, and at the 2nd one, $x^{i}$ is transformed into $y^{i}$, by setting $y^{i}_{w}:=C_{w}(x^{i}_{w})$ for each $w\in W$. Then the boundary function is updated according to rule (3.1). Let $F^{i=}$ denote the set of fully filled vertices $f$ in $F$ for $x^{i}$ (i.e., satisfying $|x^{i}_{f}|=q(f)$), and $W^{i=}$ the set of fully filled vertices in $W$ for $y^{i}$. For $f\in F$, let $L^{i}_{f}$ denote the set of edges $e\in E_{f}$ such that $b^{i}(e)<b(e)$. We observe the following.
(i) For each $f\in F$, the sets $L_{f}$ are monotone non-decreasing: $L^{i}_{f}\subseteq L^{i+1}_{f}$, and the current value on each edge $e\in L_{f}$ may decrease (at the 2nd stages of some iterations) but never increases (since $b^{1}(e)\geq\cdots\geq b^{i}(e)\geq b^{i+1}(e)\geq\cdots$).
@@
(iii) The heads $H_{w}(y)$ of vertices $w\in W^{\bullet=}$ are monotone non-decreasing: $H_{w}(y^{i})\subseteq H_{w}(y^{i+1})$ (when $w\in W^{i=}$), unless the critical tie in $\Pi_{w}$ changes (becomes better). In turn, the heads $H_{j}(x)$ for $f\in F^{\bullet=}$ are monotone non-increasing: $H_{f}(x^{i})\supseteq H_{f}(x^{i+1})$.
@@
(iv) For each $f\in F$, all edges in $E_{f}-(H^{i}_{f}\cup L^{i}_{f})$ are saturated, where $H^{i}_{f}$ denotes the head $H_{f}(x^{i})$ (which is nonempty if $f\in F^{i=}$, and empty if $f\in F-F^{i=}$).
@@
These observations prompt monotone parameters, and we say that $i$-th iteration is positive if at least one of the following events happens in it: for some $f\in F$, the set of edges $e\in E_{f}$ that are either saturated or contained in $L_{f}$ increases; the set $W^{\bullet=}$ of fully filled vertices in $W$ increases; for some $w\in W^{\bullet=}$, either the head $H_{w}$ increases, or the critical tie for $w$ becomes better. The number of positive iterations does not exceed $2|E|+|W|$.
@@
Next we deal with a sequence $Q$ of consecutive non-positive iterations and consider a current iteration among these. At the 1st stage of this iteration, the sum of values of the current function $x$ increases (compared with the preceding $y$) by some amount $\xi>0$. Note that if for a vertex $w\in W^{\bullet=}$, the value increases on some edge $e=fw$, then this edge cannot belong to the head $H_{w}$ or to $E^{\rm end}_{w}$. (For otherwise the cutting operation at the 2nd stage of the current iteration decreases the value on $e$, and as a consequence, $e$ is added as a new element to the set $L_{f}$, which is impossible since the iteration is non-positive.)
@@
It follows that an increase on some edge $e=fw$ at the 1st stage of an iteration from $Q$ is possible only in two cases: either $w\in W^{\bullet=}$ and $e$ is contained in the current tail $T_{w}$, or $w$ belongs to the current deficit set $W-W^{\bullet=}$. This implies that the total increase $\xi^{\prime}$ on the edges incident to vertices in $W^{\bullet=}$ satisfies $\xi^{\prime}\leq\xi$ (the equality may happen only if there is no contribution from $\xi$ to the vertices in $W-W^{\bullet=}$). At the 2nd stage of this iteration, the total value in the heads $H_{w}$ for these vertices $w\in W^{\bullet=}$ must decrease by the same amount $\xi^{\prime}$ (emphasize that these heads $H_{w}$ do not change, and each edge $fw$ where the value decreases has been added to the set $L_{f}$ on earlier iterations). Then we come to the 1st stage of the next iteration with the total “quota-underloading” $\xi^{\prime}$ in some vertices in $F$, which further causes a “quota-overloading” $\xi^{\prime\prime}\leq\xi^{\prime}$ in $W^{\bullet=}$, and so on. When such “overloadings” decrease geometrically, non-positive iterations may continue infinitely long (yet providing a convergence to a limit function).
@@
Relying on the observations above, we are now going to aggregate iterations from $Q$. More precisely, consider the first iteration in $Q$; assume that it is $i$-th iteration in the ordinary process. For brevity, we denote $x^{i}$ by $x$, and $y^{i}$ by $y$ (the functions obtained on the 1st and 2nd stages of the first iteration in $Q$). Also denote:
@@
The sequence $Q$ without the first iteration in it is replaced by one “big” (aggregated) iteration. To do so, we arrange the following linear program with variables $\phi_{f}$ for $f\in F^{=}$, and $\psi_{w}$ for $w\in W^{=}$:
@@
Define $y^{\prime}:=y+\Delta^{\phi,\psi}$. Then, based on the properties of $y$, we observe the following: (A.2) implies that all vertices in $W^{=}$ are fully filled for $y^{\prime}$; (A.3) implies that $y^{\prime}$ is nonnegative; (A.4) implies that $y^{\prime}$ does not violate the quotas within $F^{=}$; (A.5) implies that $y^{\prime}$ does not exceed the upper bounds $b$; (A.6) can be rewritten as $y^{\prime}(e)=a(y,w)-\psi_{w}\geq y(fw)+\phi_{f}=y^{\prime}(fw)$, where $e\in H_{w}$, implying that the edges in the head $H_{w}$ for $y$ belong to the head at $w$ for $y^{\prime}$ as well; and (A.7) implies that $y^{\prime}$ does not violate the quotas within $W-W^{=}$.
@@
Suppose that none of the constraints in (A.3),(A.5)–(A.7) turns into equality, and therefore, the attained constraints concern only (A.4). Then the resulting function $y^{\prime}$ is well-defined, in the sense that it is suitable to applying further ordinary iterations. Let $x^{\prime\prime}$ be the function obtained by applying to $y^{\prime}$ the 1st stage of the ordinary iteration. We observe that if ${\cal L}$ involves all inequalities in (A.4), then all vertices in $F^{=}$ are fully filled for $y^{\prime}$; this implies that $x^{\prime\prime}=y^{\prime}$, and the process terminates.
Now suppose that there is an inequality in (A.4) that is strict for $y^{\prime}$, and therefore, the subset $F^{\prime}$ formed by $f\in F^{=}$ with $|y^{\prime}_{f}|<q(f)$ is nonempty. Let $y^{\prime\prime}$ be the functions obtained by applying to $x^{\prime\prime}$ the 2nd stage of the ordinary iteration. Then $\delta:=y^{\prime\prime}-y^{\prime}$ is represented as $\Delta^{\phi^{\prime},\psi^{\prime}}$, where $\phi^{\prime}$ is a function on $F^{=}$ such that $\phi^{\prime}_{f}>0$ for each $f\in F^{\prime}$. It follows that the pair $(\phi^{\prime\prime},\psi^{\prime\prime}):=(\phi+\phi^{\prime},\psi+\psi^{\prime})$ gives a feasible solution to (A.1)–(A.8). But $\eta(\phi^{\prime\prime})=\eta(\phi)+\eta(\phi^{\prime})>\eta(\phi)$, contrary to the optimality of $(\phi,\psi)$.
Thus, the big iteration either terminates the whole process or changes at least one of the above parameters. Also in the modified method as described above, each big iteration is preceded by one ordinary iteration and there are no two consecutive ordinary non-positive iterations. Therefore, the whole process finishes after $O(|E|)$ iterations. Each big iteration is performed in time polynomial in $|E|$, by using a strongly polynomial l.p. method (see, e.g., [21]) to solve program (A.1)–(A.8).
@@
It remains to ask: whether the resulting $\widehat{x}$ found by the modified algorithm does coincide with the $F$-optimal assignment $x^{\rm min}$? In essence this question is not so important to us, because once we are given one or another stable assignment, we can always transform it into $x^{\rm min}$ by use of a sequence of $O(|E|)$ rotational shifts, by considering the polar order $\succeq_{W}$ on ${\cal S}$ and acting as in Sect. 6. As a result, we can conclude with the following
@@
The above modified algorithm to find $x^{\rm min}$ is finite and consists of $O(|E|)$ iterations, each taking time polynomial in $|E|$. $\Box$
Remark 6.  As a (partial) alternative to the above method, we can offer an efficient and relatively simple “combinatorial” algorithm which, given an instance of SMP, decides whether a stable solution to it is quota filling (i.e. any stable solution attains the quotas for all vertices), and if does, finds a stable solution. To do so, the input graph $G=(F\sqcup W,E)$ is extended to graph $G^{\prime}=(F^{\prime}\sqcup W^{\prime},E^{\prime})$ by adding:
vertex $f_{0}$ (to $F$) and vertex $w_{0}$ (to $W$);
edges $f_{0}w$ for all $w\in W$, denoting the set of these edges by $A$;
edges $fw_{0}$ for all $f\in F$, denoting the set of these edges by $B$;
edge $f_{0}w_{0}$.
The capacities $b$ are extended to $E^{\prime}-E$ as
  1. (A.9)
    $b(f_{0}w):=q(w)$, $w\in W$;   $b(fw_{0}):=q(f)$, $f\in F$;   and $b(f_{0}w_{0}):=\infty$,
and the quotas are extended as
  1. (A.10)
    $q(f_{0}):=\sum(q(w)\colon w\in W)=:Q_{W}$  and  $q(w_{0}):=\sum(q(f)\colon f\in F=:Q_{F}$.
The preferences are extended as follows (keeping the old orders $\geq_{v}$ within $G$):
  1. (A.11)
    • (a)
      for $f\in F$ and all $e\in E_{f}$, assign $fw_{0}\succ_{f}e$;
    • (b)
      for $w\in W$ and all $e\in E_{w}$, assign $e\succ_{w}f_{0}w$;
    • (c)
      assign $e\succ_{f_{0}}f_{0}w_{0}$ for all $e\in A$,  and assign $f_{0}w_{0}\succ_{w_{0}}e^{\prime}$ for all $e^{\prime}\in B$;
    • (d)
      assign arbitrary preferences on $A$ w.r.t. $f_{0}$, and on $B$ w.r.t. $w_{0}$.
For convenience we will use notation with symbol $y$ for assignments on $E^{\prime}$, and with $x$ for the restrictions of $y$ on $E$. Define $y_{0}$ on $E^{\prime}$ as
  1. (A.12)
    $y_{0}(e):=b(e)$ for all $e\in A$,  $y_{0}(e^{\prime}):=b(e^{\prime})$ for all $w\in W$, and 0 otherwise.
This $y_{0}$ is quota filling (in $G^{\prime}$) and stable. Moreover ((A.11))(a) ensures that $y_{0}$ is the best assignment for $F^{\prime}$, i.e. $y_{0}=y^{\rm min}$. One can show the following
Lemma A.2
(i) Let $x$ be a stable assignment for $G$ and suppose that $x$ is quota filling (which implies that $Q_{F}=Q_{W}$). Let $y$ be the extension of $x$ to $E^{\prime}$ that takes zero values on $A\cup B$ and satisfies $y(f_{0}w_{0})=Q_{F}$. Then $y$ is a stable assignment for $G^{\prime}$.
[…]
Proof  (i) follows from the facts that: $y$ is quota filling (since $y(f_{0}w_{0})=q(f_{0})=q(w_{0})$); and no edge in $E^{\prime}-E=A\cup B\cup\{f_{0}w_{0}\}$ is blocking for $y$ (since (a),(b),(c) in ((A.11)) imply that: for each $w\in W$, $f_{0}w$ is the worst in $E^{\prime}_{w}$; for each $f\in F$, $fw_{0}$ is worse than the unique edge $f_{0}w_{0}$ in $E^{\prime}_{w_{0}}$ where $y$ is nonzero; and $f_{0}w_{0}$ is the worst in $E^{\prime}_{f_{0}}$).
Now we show (ii). Observe that any $w\in W$ is fully filled by $x$. For otherwise $w$ is deficit for $y$ (since $y(E^{\prime}_{w})=x(E_{w})+y(f_{0}w)=x(E_{w})<q(w)$), implying that the edge $e=f_{0}w$ is blocking for $y$ in both cases: $y(f_{0}w_{0})<q(f_{0})$, and $y(f_{0}w_{0})=q(f_{0})$ (in view of $e\succ_{f_{0}}f_{0}w_{0}$ and $q(f_{0})=Q_{W}>0$). So we have $|x_{w}|=q(w)$ for all $w\in W$, and therefore,
$Q_{W}=\sum(x(E_{w})\colon w\in W)=\sum(x(E_{f})\colon f\in F)\leq Q_{F}.$
The case $Q_{W}<Q_{F}$ is impossible. For otherwise we would have $y(E^{\prime}_{w_{0}})=y(f_{0}w_{0})<q(w_{0})=Q_{F}$ and $y(E^{\prime}_{f})=x(E_{f})<q(f)$ for some $f\in F$, and therefore, the edge $fw_{0}$ would be blocking for $y$ (as it is unsaturated and connects two deficit vertices). Thus, $Q_{W}=Q_{F}$. This implies that $x$ is quota filling. Now the stability of $x$ easily follows. $\Box$
Based on Lemma A.2, we apply to $G^{\prime}$ the rotational method to obtain $y^{\rm max}$, by drawing a route from $y^{\rm min}$ to $y^{\rm max}$. Let $x$ be the restriction of $y=y^{\rm max}$ to $E$. If $y$ takes zero values on $A\cup B$, then, by (ii) in the lemma, $x$ is stable and quota filling. This implies that $x=x^{\rm max}$ (since $x^{\rm max}$ admits a proper extension to $G^{\prime}$, according to (i) in the lemma). And if $y$ is not identically zero on $A\cup B$, then we can declare that the stable assignments for $G$ are not quota filling; this again follows from (i) in the lemma.
Thus, we obtain an efficient algorithm, not appealing to linear programming, to recognize the existence of quota filling assignments for a given instance of SMP and find one of them. If desirable, we can further efficiently transform $x^{\rm max}$ into $x^{\rm min}$, by using “reversed” rotations in $G$.
[…]